# 1]organization=IIT-CNR, addressline=Via Moruzzi 1, code=56124, city=Pisa, country=Italy

### v3

@@ 4 Data Collection and Preparation
For this study we use the dataset from [11, 18], consisting of 16 million Reddit comments made by 16,828 different users that were affected by the Great Ban.
@@ 4.1 Banned subreddits selection
Each user in the dataset was an active participant before the ban in at least one of 15 selected subreddits. The selection of the 15 subreddits was done as follows. Out of the 2,000 subreddits shut during the Great Ban, Reddit only publicly disclosed the names of the 10 largest ones. [^1] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 03/15/2024) For all remaining banned subreddits, Reddit only provided a list of partially obfuscated names. M. Trujillo et al. [21] deciphered the list, removed all private subreddits as well as all those having less than 2,000 active users. Only 5 subreddits remained from the obfuscated list, which M. Trujillo et al. [21] and Cima et al. [11] studied in addition to the 10 publicly disclosed ones. Table 1 provides the list of the 15 subreddits on which our study is focused, together with some descriptive statistics.
@@ 4.2 Active users selection
For each of the 15 selected subreddits, Cima et al. [11] gathered all comments posted therein between December 2019 and June 2020, which resulted in a total of 8 million comments posted by approximately 194K distinct users. As shown in Figure 3, the timeframe used for the data collection spans 7 months leading up to the Great Ban, providing a strong reference for the activity levels of the affected users before the moderation intervention [14]. The initial set of 194K users was later filtered so as to retain only those users who showed consistent activity in at least one of the 15 selected subreddits. In detail, only those users who posted at least one comment each month were retained. In [11], this filtering step was useful to obtain meaningful post-hoc estimations of the effect of the Great Ban on the activity of the users. Since here we also study changes in user activity, albeit in a predictive fashion, this filtering step is suitable for our analysis as well. In addition, bots were also removed from the dataset. First, we followed reference practices in literature and we removed all accounts that posted two or more comments in less than a second [46]. Subsequently, we manually analyzed a random sample of 1,000 removed accounts to ensure that only bots were removed from the dataset. A similar analysis also revealed that among the remaining users in the dataset there were no clearly distinguishable bots. As a result of these filtering and validation steps, we obtained the dataset DB-B composed of 2.2 million comments posted within the 15 selected subreddits by 16,828 distinct users, as summarized in Table 1.
@@ 4.3 Data from outside the banned subreddits
Evaluating (or predicting) the effect of a moderation intervention involves comparing data from before and after the intervention itself. This requirement also surfaces from the definitions that we gave in Section 3. However, no activity exists post-intervention within the banned subreddits, since they were all permanently shut. Therefore, data about the behavior of the affected users outside of the banned subreddits must be used. For this reason, Cima et al. [11] collected all comments made by the 16,828 selected users outside of the 15 banned subreddits over a wide timeframe spanning 7 months before and 7 months after the Great Ban, as depicted in Figure 3. This additional data collection yielded approximately 13.8 million comments, of which 8.2 million (59%) were posted before the ban and constitute dataset DN-B while 5.6 million (41%) were posted afterwards and constitute dataset DN-A. We used the data from before the ban (DN-B) to compute our machine learning features, while the data posted after the ban (DN-A) allows assigning ground-truth labels to the users based on their post-ban activity. Notably, out of the 16,828 initially selected users, 288 (1.7%) did not have any activity outside of the 15 banned subreddits and were thus discarded.
@@ 5 Machine Learning Approach
We begin by extracting user-level features from the portion of the dataset related to the activity of the selected users before the ban, both within (DB-B) and outside (DN-B) of the banned subreddits. Before further processing this data we perform outliers removal, as this helps enhancing the robustness and generalization of the trained models by mitigating the potential distortions and biases caused by extreme data points [47]. Data about user activities outside of the banned subreddits after the ban (DN-A) is used to assign ground-truth labels to the users. User-level features are then merged with the ground-truth labels and the resulting dataset is split into a training and a test set. The test set is only used for the final evaluation of the optimized models. Instead, the training set is used for feature selection and for optimizing and training the models. Given that we experiment with multiple classification models, each with its own characteristics and hyperparameters, the feature selection and optimization steps of our machine learning pipeline ensure that each model operates in optimal conditions. In other words, our approach ensures meaningful and fair comparisons between the different models. The first step in the model optimization process involves selecting an adequate number and set of features for each model. Specifically, we rank all features and we select the top-N ones to reduce the dimensionality of our models and to eliminate redundant and correlated features. For each model we experiment with multiple numbers N of selected features, as different models might benefit from a smaller or larger number of features. To have accurate and robust estimates of the model performances with the different sets of features, we carry out a 10-fold cross-validation on the training set. At each iteration, we rescale features so that they all cover a similar range of values. Rescaling ensures that no feature dominates the learning process due to overly large magnitude, and also contributes to speeding up training times  [48, 49]. At each iteration we also rebalance the training set. In fact, both the hard and soft abandonment tasks are heavily imbalanced, given that abandoning users constituted a small minority of all users (i.e., 14.9% for hard abandonment and 26.9% for soft abandonment). We note that this imbalance is common for tasks concerning the prediction of human behavior [50]. Therefore, by rebalancing the training set we reduce the bias that the trained models might exhibit towards those users who did not abandon the platform (i.e., the majority class). The averaged results of the 10-fold cross-validation process are used to select the optimal number and set of features for each model. Up to now, all models were trained and evaluated with default hyperparameters. In the last step of model optimization, for each model we perform a grid search with cross-validation over its hyperparameters so as to choose the best combination of them. Finally, we use the selected set of features and hyperparameters to train an optimized version of each model on the whole training set. Feature scaling and rebalancing are applied also for hyperparameter tuning and model training. Each optimized model is then evaluated on the held-out test set. Evaluation results and comparisons are presented in Section 8.
@@ 6.1 Activity features
Features in this class describe the overall level of engagement and participation of a user within Reddit. The motivation for including activity features is straightforward, given our goal of predicting user abandonment—that is, the interruption of posting activity on the platform. Moreover, previous descriptive studies on the effects of moderation interventions, such as those reviewed in Section 2.1, showed that bans frequently cause a reduction in user activity [11, 21, 13, 18], which supports the adoption of this type of features in our task. Some examples of the activity features that we implemented are the total number of comments posted by a user, the average time between the sharing of two subsequent comments, the slope of the trend of posted comments in either the banned and the non-banned subreddits, which captures whether user activity in certain subreddits was increasing or decreasing prior to the ban. Other features of this class capture possible peculiar characteristics of the comments, one of them being the number of stickied comments. This feature was included as it was among the strongest predictors of ban evasion in a previous study [31].
@@ 6.2 Toxicity features
The users considered in our study were active participants in subreddits banned due to the widespread presence of toxic content. As such, features gauging the degree of toxicity of the users pre-ban could be strong predictors in our task. In addition, integrating toxicity features into the classification models allows exploring the extent to which differences in toxic behavior are capable of explaining user reactions to the ban. We computed toxicity scores with Detoxify [51], a state-of-the-art [11] multilingual deep learning toxicity classifier that is widely used in literature [52, 53], including predictive studies for the detection of cyberbullying [54, 55]. Given a piece of text, Detoxify provides multiple indicators of toxicity, such as the toxicity and severe toxicity scores, as well as additional scores for obscenity, insults, identity attacks, and threats. In addition, we also computed sentiment scores given the strong relationship between toxicity and sentiment [56]. For obtaining sentiment scores we used Vader, a well-known rule-based sentiment analyzer that is specifically designed for social media content [57]. Vader outputs positive, negative, neutral, and compound sentiment scores, with the latter being the sum of the previous ones. To compute toxicity and sentiment user-level features we first classified each comment by each user with Detoxify and Vader. We then aggregated each score provided by the two tools over all comments by the same user, by computing the mean, minimum, maximum, and standard deviation of the scores, obtaining the $4\times 10=40$ features in this class.
@@ 6.3 Writing style features
Several previous works demonstrated that users may exhibit linguistic changes after suffering a moderation intervention. For example, [7] noted changes in the use of the first and third plural pronouns, as well as in word choices, after a ban. Similarly, [21] found that users who remained on Reddit after a ban exhibited reduced use of in-group language. Other prior research has shown that moderation interventions can lead to linguistic shifts, including changes in word choices, sentiment, and rhetorical strategies. Studies have observed that moderated users may adjust their language to evade detection, adopt more aggressive or explicit speech in response to relaxed moderation policies, or even develop lexical variations of restricted terms to bypass enforcement mechanisms. These findings suggest that subtle linguistic cues may reflect user adaptation, disengagement, or defiance following an intervention [58, 21, 59]. Based on these preliminary descriptive results, here we assess the extent to which basic linguistic and writing style features provide predictive information about the activity of the users after a ban. Among the features that we computed are counts of the different parts-of-speech used and readability scores such as the well-known Flesch-Kincaid Grade Level and the SMOG Index. Each of these scores were computed for each user comment and subsequently aggregated at the user level by computing the mean, minimum, maximum, and standard deviation.
@@ 6.4 Relational features
The relationships between users and communities constitute the fundamental fabric upon which social media platforms are woven, shaping the dynamics of online engagement and interaction. Moreover, changes in user-user and user-subreddit relationships have already been observed as a consequence of moderation interventions [13, 31]. For this reason, considering relational features in our task allows us to capture the interaction dynamics that are inherent to Reddit communities. As an example, we computed features that quantify the degree of influence that each user had in the banned subreddits, which could be predictive of future abandonment. This feature was computed by ranking users based on the average score (i.e., the difference between upvotes and downvotes) of their comments in the banned subreddits. In addition, the same feature was also computed for the non-banned subreddits. Similarly, we examined the relationship between users and specific subreddits, identifying those users who predominantly participated in the banned subreddits, which could provide an early sign of abandonment. In fact, users who almost exclusively participated in banned subreddits could lack motivation to stay on the platform after those subreddits were shut [7]. Finally, we leveraged results of recent works that demonstrated the usefulness of initiative and adaptability features in social media user classification tasks [60]. For example, we measured the number and ratio of threads started by each user as a proxy for their capacity to drive the conversation in a subreddit, rather than to follow what others say.
@@ 7.1 Comparisons
We experiment with multiple reference classification models to solve the hard and soft abandonment tasks, including Naive Bayes (NB), K-Nearest Neighbors (KNN), Decision Tree (DT), Random Forest (RF), Adaptive Boosting (AB), Gradient Boosting (GB), and Support Vector Machine (SVM). This selection spans the main families of classification algorithms, encompassing instance-based, probabilistic, ensemble-based, and discriminative methods of different complexities, thus providing an extensive exploration of many learning paradigms. Furthermore, similar choices of classification models were made in recent works on related tasks, such as the detection of different forms of online harms [61, 62, 63, 64]. In addition to the previous classification models, we also implement five simple baselines for further comparison. The Stratified baseline generates predictions by following the ground-truth label distribution, without taking any feature into account. The DT Ratio baseline employs a decision tree that receives a single feature in input representing the ratio of comments made by a user in the non-banned subreddits over the comments made by that user in the banned ones. The rationale for this baseline is that users who are more active on non-banned subreddits are less prone to abandon the platform. The DT Trend baseline employs a decision tree that receives a single feature in input representing the trend of the number of monthly comments by a user before the ban. The reason for including this baseline is that users with a decreasing posting trend are more likely to abandon the platform. Finally, to test the effectiveness of our machine learning pipeline, we train a Naive Bayes (NB-NFE) and a Decision Tree (DT-NFE) model by only performing feature standardization, without any other step of our pipeline —namely, rebalancing, feature selection, and hyperparameter tuning. For the sake of simplicity, we do not apply any preprocessing steps to the baselines. Finally, we remark that we resort to using these baselines instead of more powerful approaches because the novelty of our task makes it so that no previous work exists on predicting the effects of a moderation intervention.
@@ 7.2.1 Outliers removal
To carry out outliers detection we apply Isolation Forest, a decision tree-based method that is particularly suitable for identifying anomalies in highly-dimensional data, such as ours [65]. Isolation Forest constructs decision trees by randomly selecting features and split points, efficiently isolating anomalies by assigning them shorter average path lengths in the trees compared to normal data instances, by leveraging the natural tendency of anomalies to require fewer splits for isolation. The application of Isolation Forest to our dataset led to the detection and removal of 297 outliers. A manual verification revealed that the removed users featured extremely low levels of activity, which supports their removal.
@@ 7.2.4 Feature scaling and rebalancing
During the feature selection, hyperparameter tuning, and model training steps, we perform feature scaling for all models. Additionally, for all models except for the ensemble-based ones —namely, RF, AB, and GB— we also perform data rebalancing. For ensemble-based models, removing data rebalancing often improves performance, as also confirmed by our experimental results, because these models naturally handle class imbalance and can be negatively affected by data rebalancing, which might distort data distributions [66]. Feature scaling is applied first, via Z-score standardization, so that each feature has mean $\mu=0$ and standard deviation $\sigma=1$. Rebalancing is applied last, as suggested in previous literature [67]. For rebalancing, we leverage both oversampling and undersampling techniques to mitigate the large class imbalance of our dataset. We perform oversampling with SMOTE, a state-of-the-art technique that creates new artificial samples of the minority class [67]. Instead, we use Random Undersampling to randomly discard samples from the majority class [68]. For both tasks we use the two techniques in combination to reach a final fixed imbalance of 60/40 between the majority and minority class, down from 85/15 for hard abandonment and 73/27 for soft abandonment. This approach allows us to reduce the skewness of the class distribution without creating too many artificial samples that could introduce bias, and without discarding too much data that could result in degraded model performance.

### v4

@@ 4 Data Collection and Preparation
For this study we use the dataset from [cima2024great, cima2025investigating], consisting of 16 million Reddit comments made by 16,828 different users that were affected by the Great Ban.
@@ 4.1 Banned subreddits selection
Each user in the dataset was an active participant before the ban in at least one of 15 selected subreddits. The selection of the 15 subreddits was done as follows. Out of the 2,000 subreddits shut during the Great Ban, Reddit only publicly disclosed the names of the 10 largest ones. [^1] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 03/15/2024) For all remaining banned subreddits, Reddit only provided a list of partially obfuscated names. M. Trujillo et al. [trujillo2021echo] deciphered the list, removed all private subreddits as well as all those having less than 2,000 active users. Only 5 subreddits remained from the obfuscated list, which M. Trujillo et al. [trujillo2021echo] and Cima et al. [cima2024great] studied in addition to the 10 publicly disclosed ones. Table 1 provides the list of the 15 subreddits on which our study is focused, together with some descriptive statistics.
@@ 4.2 Active users selection
For each of the 15 selected subreddits, Cima et al. [cima2024great] gathered all comments posted therein between December 2019 and June 2020, which resulted in a total of 8 million comments posted by approximately 194K distinct users. As shown in Figure 3, the timeframe used for the data collection spans 7 months leading up to the Great Ban, providing a strong reference for the activity levels of the affected users before the moderation intervention [trujillo2023one]. The initial set of 194K users was later filtered so as to retain only those users who showed consistent activity in at least one of the 15 selected subreddits. In detail, only those users who posted at least one comment each month were retained. In [cima2024great], this filtering step was useful to obtain meaningful post-hoc estimations of the effect of the Great Ban on the activity of the users. Since here we also study changes in user activity, albeit in a predictive fashion, this filtering step is suitable for our analysis as well. In addition, bots were also removed from the dataset. First, we followed reference practices in literature and we removed all accounts that posted two or more comments in less than a second [hurtado2019bot]. Subsequently, we manually analyzed a random sample of 1,000 removed accounts to ensure that only bots were removed from the dataset. A similar analysis also revealed that among the remaining users in the dataset there were no clearly distinguishable bots. As a result of these filtering and validation steps, we obtained the dataset DB-B composed of 2.2 million comments posted within the 15 selected subreddits by 16,828 distinct users, as summarized in Table 1.
@@ 4.3 Data from outside the banned subreddits
Evaluating (or predicting) the effect of a moderation intervention involves comparing data from before and after the intervention itself. This requirement also surfaces from the definitions that we gave in Section 3. However, no activity exists post-intervention within the banned subreddits, since they were all permanently shut. Therefore, data about the behavior of the affected users outside of the banned subreddits must be used. For this reason, Cima et al. [cima2024great] collected all comments made by the 16,828 selected users outside of the 15 banned subreddits over a wide timeframe spanning 7 months before and 7 months after the Great Ban, as depicted in Figure 3. This additional data collection yielded approximately 13.8 million comments, of which 8.2 million (59%) were posted before the ban and constitute dataset DN-B while 5.6 million (41%) were posted afterwards and constitute dataset DN-A. We used the data from before the ban (DN-B) to compute our machine learning features, while the data posted after the ban (DN-A) allows assigning ground-truth labels to the users based on their post-ban activity. Notably, out of the 16,828 initially selected users, 288 (1.7%) did not have any activity outside of the 15 banned subreddits and were thus discarded.
@@ 5 Machine Learning Approach
We begin by extracting user-level features from the portion of the dataset related to the activity of the selected users before the ban, both within (DB-B) and outside (DN-B) of the banned subreddits. Before further processing this data we perform outliers removal, as this helps enhancing the robustness and generalization of the trained models by mitigating the potential distortions and biases caused by extreme data points [smiti2020critical]. Data about user activities outside of the banned subreddits after the ban (DN-A) is used to assign ground-truth labels to the users. User-level features are then merged with the ground-truth labels and the resulting dataset is split into a training and a test set. The test set is only used for the final evaluation of the optimized models. Instead, the training set is used for feature selection and for optimizing and training the models. Given that we experiment with multiple classification models, each with its own characteristics and hyperparameters, the feature selection and optimization steps of our machine learning pipeline ensure that each model operates in optimal conditions. In other words, our approach ensures meaningful and fair comparisons between the different models. The first step in the model optimization process involves selecting an adequate number and set of features for each model. Specifically, we rank all features and we select the top-N ones to reduce the dimensionality of our models and to eliminate redundant and correlated features. For each model we experiment with multiple numbers N of selected features, as different models might benefit from a smaller or larger number of features. To have accurate and robust estimates of the model performances with the different sets of features, we carry out a 10-fold cross-validation on the training set. At each iteration, we rescale features so that they all cover a similar range of values. Rescaling ensures that no feature dominates the learning process due to overly large magnitude, and also contributes to speeding up training times  [ozsahin2022impact, wan2019influence]. At each iteration we also rebalance the training set. In fact, both the hard and soft abandonment tasks are heavily imbalanced, given that abandoning users constituted a small minority of all users (i.e., 14.9% for hard abandonment and 26.9% for soft abandonment). We note that this imbalance is common for tasks concerning the prediction of human behavior [robertson2022uncommon]. Therefore, by rebalancing the training set we reduce the bias that the trained models might exhibit towards those users who did not abandon the platform (i.e., the majority class). The averaged results of the 10-fold cross-validation process are used to select the optimal number and set of features for each model. Up to now, all models were trained and evaluated with default hyperparameters. In the last step of model optimization, for each model we perform a grid search with cross-validation over its hyperparameters so as to choose the best combination of them. Finally, we use the selected set of features and hyperparameters to train an optimized version of each model on the whole training set. Feature scaling and rebalancing are applied also for hyperparameter tuning and model training. Each optimized model is then evaluated on the held-out test set. Evaluation results and comparisons are presented in Section 8.
@@ 6.1 Activity features
Features in this class describe the overall level of engagement and participation of a user within Reddit. The motivation for including activity features is straightforward, given our goal of predicting user abandonment—that is, the interruption of posting activity on the platform. Moreover, previous descriptive studies on the effects of moderation interventions, such as those reviewed in Section 2.1, showed that bans frequently cause a reduction in user activity [cima2024great, trujillo2021echo, trujillo2022make, cima2025investigating], which supports the adoption of this type of features in our task. Some examples of the activity features that we implemented are the total number of comments posted by a user, the average time between the sharing of two subsequent comments, the slope of the trend of posted comments in either the banned and the non-banned subreddits, which captures whether user activity in certain subreddits was increasing or decreasing prior to the ban. Other features of this class capture possible peculiar characteristics of the comments, one of them being the number of stickied comments. This feature was included as it was among the strongest predictors of ban evasion in a previous study [habib2022proactive].
@@ 6.2 Toxicity features
The users considered in our study were active participants in subreddits banned due to the widespread presence of toxic content. As such, features gauging the degree of toxicity of the users pre-ban could be strong predictors in our task. In addition, integrating toxicity features into the classification models allows exploring the extent to which differences in toxic behavior are capable of explaining user reactions to the ban. We computed toxicity scores with Detoxify [Detoxify], a state-of-the-art [cima2024great] multilingual deep learning toxicity classifier that is widely used in literature [bagdasaryan2022spinning, kopf2024openassistant], including predictive studies for the detection of cyberbullying [teng2023cyberbullying, ejaz2024towards]. Given a piece of text, Detoxify provides multiple indicators of toxicity, such as the toxicity and severe toxicity scores, as well as additional scores for obscenity, insults, identity attacks, and threats. In addition, we also computed sentiment scores given the strong relationship between toxicity and sentiment [gilda2022predicting]. For obtaining sentiment scores we used Vader, a well-known rule-based sentiment analyzer that is specifically designed for social media content [vader]. Vader outputs positive, negative, neutral, and compound sentiment scores, with the latter being the sum of the previous ones. To compute toxicity and sentiment user-level features we first classified each comment by each user with Detoxify and Vader. We then aggregated each score provided by the two tools over all comments by the same user, by computing the mean, minimum, maximum, and standard deviation of the scores, obtaining the $4\times 10=40$ features in this class.
@@ 6.3 Writing style features
Several previous works demonstrated that users may exhibit linguistic changes after suffering a moderation intervention. For example, [horta2021platform] noted changes in the use of the first and third plural pronouns, as well as in word choices, after a ban. Similarly, [trujillo2021echo] found that users who remained on Reddit after a ban exhibited reduced use of in-group language. Other prior research has shown that moderation interventions can lead to linguistic shifts, including changes in word choices, sentiment, and rhetorical strategies. Studies have observed that moderated users may adjust their language to evade detection, adopt more aggressive or explicit speech in response to relaxed moderation policies, or even develop lexical variations of restricted terms to bypass enforcement mechanisms. These findings suggest that subtle linguistic cues may reflect user adaptation, disengagement, or defiance following an intervention [chancellor2016thyghgapp, trujillo2021echo, wadden2021effect]. Based on these preliminary descriptive results, here we assess the extent to which basic linguistic and writing style features provide predictive information about the activity of the users after a ban. Among the features that we computed are counts of the different parts-of-speech used and readability scores such as the well-known Flesch-Kincaid Grade Level and the SMOG Index. Each of these scores were computed for each user comment and subsequently aggregated at the user level by computing the mean, minimum, maximum, and standard deviation.
@@ 6.4 Relational features
The relationships between users and communities constitute the fundamental fabric upon which social media platforms are woven, shaping the dynamics of online engagement and interaction. Moreover, changes in user-user and user-subreddit relationships have already been observed as a consequence of moderation interventions [trujillo2022make, habib2022proactive]. For this reason, considering relational features in our task allows us to capture the interaction dynamics that are inherent to Reddit communities. As an example, we computed features that quantify the degree of influence that each user had in the banned subreddits, which could be predictive of future abandonment. This feature was computed by ranking users based on the average score (i.e., the difference between upvotes and downvotes) of their comments in the banned subreddits. In addition, the same feature was also computed for the non-banned subreddits. Similarly, we examined the relationship between users and specific subreddits, identifying those users who predominantly participated in the banned subreddits, which could provide an early sign of abandonment. In fact, users who almost exclusively participated in banned subreddits could lack motivation to stay on the platform after those subreddits were shut [horta2021platform]. Finally, we leveraged results of recent works that demonstrated the usefulness of initiative and adaptability features in social media user classification tasks [mazza2022investigating]. For example, we measured the number and ratio of threads started by each user as a proxy for their capacity to drive the conversation in a subreddit, rather than to follow what others say.
@@ 7.1 Comparisons
We experiment with multiple reference classification models to solve the hard and soft abandonment tasks, including Naive Bayes (NB), K-Nearest Neighbors (KNN), Decision Tree (DT), Random Forest (RF), Adaptive Boosting (AB), Gradient Boosting (GB), and Support Vector Machine (SVM). This selection spans the main families of classification algorithms, encompassing instance-based, probabilistic, ensemble-based, and discriminative methods of different complexities, thus providing an extensive exploration of many learning paradigms. Furthermore, similar choices of classification models were made in recent works on related tasks, such as the detection of different forms of online harms [cresci2015fame, liu2016detecting, tardelli2022detecting, amira2023detection]. In addition to the previous classification models, we also implement five simple baselines for further comparison. The Stratified baseline generates predictions by following the ground-truth label distribution, without taking any feature into account. The DT Ratio baseline employs a decision tree that receives a single feature in input representing the ratio of comments made by a user in the non-banned subreddits over the comments made by that user in the banned ones. The rationale for this baseline is that users who are more active on non-banned subreddits are less prone to abandon the platform. The DT Trend baseline employs a decision tree that receives a single feature in input representing the trend of the number of monthly comments by a user before the ban. The reason for including this baseline is that users with a decreasing posting trend are more likely to abandon the platform. Finally, to test the effectiveness of our machine learning pipeline, we train a Naive Bayes (NB-NFE) and a Decision Tree (DT-NFE) model by only performing feature standardization, without any other step of our pipeline —namely, rebalancing, feature selection, and hyperparameter tuning. For the sake of simplicity, we do not apply any preprocessing steps to the baselines. Finally, we remark that we resort to using these baselines instead of more powerful approaches because the novelty of our task makes it so that no previous work exists on predicting the effects of a moderation intervention.
@@ 7.2.1 Outliers removal
To carry out outliers detection we apply Isolation Forest, a decision tree-based method that is particularly suitable for identifying anomalies in highly-dimensional data, such as ours [liu2008isolation]. Isolation Forest constructs decision trees by randomly selecting features and split points, efficiently isolating anomalies by assigning them shorter average path lengths in the trees compared to normal data instances, by leveraging the natural tendency of anomalies to require fewer splits for isolation. The application of Isolation Forest to our dataset led to the detection and removal of 297 outliers. A manual verification revealed that the removed users featured extremely low levels of activity, which supports their removal.
@@ 7.2.4 Feature scaling and rebalancing
During the feature selection, hyperparameter tuning, and model training steps, we perform feature scaling for all models. Additionally, for all models except for the ensemble-based ones —namely, RF, AB, and GB— we also perform data rebalancing. For ensemble-based models, removing data rebalancing often improves performance, as also confirmed by our experimental results, because these models naturally handle class imbalance and can be negatively affected by data rebalancing, which might distort data distributions [khan2024review]. Feature scaling is applied first, via Z-score standardization, so that each feature has mean $\mu=0$ and standard deviation $\sigma=1$. Rebalancing is applied last, as suggested in previous literature [smote]. For rebalancing, we leverage both oversampling and undersampling techniques to mitigate the large class imbalance of our dataset. We perform oversampling with SMOTE, a state-of-the-art technique that creates new artificial samples of the minority class [smote]. Instead, we use Random Undersampling to randomly discard samples from the majority class [hasanin2018effects]. For both tasks we use the two techniques in combination to reach a final fixed imbalance of 60/40 between the majority and minority class, down from 85/15 for hard abandonment and 73/27 for soft abandonment. This approach allows us to reduce the skewness of the class distribution without creating too many artificial samples that could introduce bias, and without discarding too much data that could result in degraded model performance.